charles is 3 years younger than evan. the sun of their ages is 63, how old are they both
Found 2 solutions by ikleyn, greenestamps:
Answer by ikleyn(52781)

.

    C + (C+3) = 63


    2C + 3 = 63

    2C     = 63 - 3

    2C     = 60

     C     = 60/2 = 30.


Answer.  C is 30 years old.  E is 30 + 3 = 33 years old.

Solved.



Answer by greenestamps(13200)



I will respond to this question only to comment on how sloppy the average person is with the English language....

With the problem as posted, there is no solution because the given information is self-contradictory.

"How old are they both", according to proper English, means the two of them are the same age.

But the problem says one of them is 3 years older than the other....